WAEC Mathematics 1991 (August) Question 40

Question 40

Probability

A box contains 2 white and 3 blue identical marbles. If two marbles are picked at random, one after the other without replacement, what is the probability of picking two marbles of different colours?

  1. A.2/3
  2. B.3/5
  3. C.2/5
  4. D.7/20
  5. E.3/10
Show the answer and explanation

Answer: B (3/5)

Explanation

P(white then blue) + P(blue then white) = 2/5 × 3/4 + 3/5 × 2/4 = 6/20 + 6/20 = 3/5.